1.      One of the undersecretary of the Department of Labor and Employment (DOLE) claims that the average salary of Social worker is Php18, 000. A sample of 19 social worker’s salary has a mean of Php17, 350 and standard deviation of Php1, 230. Is there enough evidence to reject the undersecretary’s claim at 𝛼 = 0.01?


1
Expert's answer
2022-02-08T16:51:10-0500

Solution:

"H_0:\mu=18000
\\ H_1:\mu\ne18000
\\ n=19
\\ \bar X=17350
\\\sigma=1230"

Test statistics: "\dfrac{\bar X-\mu}{\sigma/\sqrt n}= \dfrac{17350-18000}{1230/\sqrt {19}}=-2.303"

"t_{n-1}=t_{18}=2.878"



This is two-tailed test, so t-value is "\pm2.878"

Our test statistics -2.303 lies between t-values. So, we cannot reject null hypothesis.

Thus, we conclude that the average salary of Social worker is Php18, 000.